# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### v2

@@ Abstract
While imaging techniques at macro and mesoscales have garnered substantial attention and resources, microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure. In this paper, we address a significant gap in the field of neuroimaging by introducing the largest-to-date public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in volume electron microscopy (VEM) images. Our BvEM benchmark is based on VEM image volumes from three mammal species: adult mouse, macaque, and human. We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation. Furthermore, we developed a zero-shot cortical blood vessel segmentation method named TriSAM, which leverages the powerful segmentation model SAM for 3D segmentation. To extend SAM from 2D to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points. This approach effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning. Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species.
Index Terms:
@@
Figure 1: Comparison of imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details. (b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their cellular details.
I Introduction
With around 2% of body weight, our brain receives around 20% of blood supply. Most of the energy and nutrients are consumed by the neurons, and neuron function is sensitive to the blood supply [3, 4]. The blood supply can even be adjusted following the consumption of brain regions, called neurovascular coupling [5]. Alterations of blood vessel structures are observed in many brain diseases, e.g., Alzheimer’s and vascular dementia [6]. Thus, blood vessels in the brain have been extensively investigated with various imaging modalities at different resolutions (Figure 1a). Compared to the macro-level imaging (e.g., CT [1] and MRI [7]) and mesoscale-level imaging (e.g., light microscopy [2]), volume electron microscopy (VEM) [8] can further reveal the detailed ultrastructure including all vascular cells (Figure 1b) for in-depth analysis. However, no large-scale annotated VEM dataset exists to develop and evaluate automated 3D blood vessel segmentation methods.
@@ I Introduction
Thus, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images to foster segmentation method development. The raw image volumes are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities. We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness. However, on the BvEM dataset, existing 3D blood vessel segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the imaging pipeline and the complexity of the 3D blood vessel morphology. Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature.
@@
To address these challenges, we propose a zero-shot 3D segmentation method, TriSAM, leveraging the recent 2D segmentation foundation model, Segment Anything Model (SAM) [15], to handle the appearance diversity. To re-purpose SAM for 3D segmentation, we developed a multi-seed segment tracking framework based upon the video object segmentation method with SAM [16]. As it is easier to track 2D blood vessel segments along the blood flow direction, TriSAM selects the best 2D plane for SAM-based tracking and further introduces a recursive seed sampling strategy that performs tri-plane selection at the potential turning points for efficiency. Experimental results demonstrate that the proposed TriSAM method significantly outperforms the prior state-of-the-art methods on the proposed BvEM benchmark across all three species.
II Related Work
@@
Blood Vessel Segmentation. Most existing VEM image segmentation algorithms were developed for neurons [17] and synapses [18, 19, 20]. The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11]. This marks a significant milestone in connectomics research. The blood vessels were segmented manually in the human cortex [12] and automatically using fully convolutional neural network (FCN) [21] or 3D U-Net [11, 22, 23]. Due to the burden of annotation, efforts have been made to decrease the need for annotations [24]. For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10]. Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters like diffusion and ridge filtering. Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
Evaluation Metrics.. Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth. The evaluation metrics could be classified into two categories: pixel-based and skeleton-based [10]. Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient [25]. Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error [26]. Skeleton-based metric requires the blood vessel skeleton that might not always be available. It also requires the computation of point-to-point correspondence between the ground truth and computed skeletons. All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets. Most of them could be utilized in VEM images directly. Here we adopt the metrics Precision, Recall, and Accuracy as we measure instance-level pixel prediction.
@@
Segment Anything-Based Models. As a foundation model for image segmentation, the recently proposed Segment Anything Model [15] has garnered significant attention and has been extended to a variety of domains [27, 28, 29, 30] including object tracking [27, 16], image inpainting [28], image mattting [31], super-resolution [32], 3D point cloud [33], and image editting [34]. Despite SAM’s remarkable generalization capabilities, it still encounters some challenges in practical applications, One of these challenges is the huge computation costs due to the heavyweight image encoder. FastSAM [35] adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed. MobileSAM [36] proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size. Therefore the MobileSAM is employed in our proposed method. One important challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical [37] or biological [38] images, particularly in the context of 3D data. Although SAM demonstrates impressive generalization capabilities, it faces challenges in specialized domains like medical [37, 39] or biological images [38], especially in handling 3D data. [30] assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties. [40] extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability. To address the domain gap between natural and medical images, SAM-Adapter [37], SAM-Med2D [41], and Medical SAM Adapter [42] introduced Adapter modules and trained the Adapter with medical images. They attained good performance on various medical image segmentation tasks. MedSAM [29] adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results. MicroSAM [38] also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data. Unlike these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
@@
III BvEM Dataset
III-A Sample Information
We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [^1] BvEM-Mouse can be visualized in: https://tinyurl.com/yc2s38x9 [11], superior temporal gyrus from an adult macaque [43], and temporal lobe from an adult human [12]. Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
TABLE I: Dataset information. Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined. The last column shows the max/total length of the blood vessel instances.
@@ III-A Sample Information
III-B Image Volume Standardization
We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for AI researchers. Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content. As shown in Fig. 2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
III-C Blood Vessel Instance Segmentation Annotation
@@
III-D Dataset Statistics
We summarize the dataset details in Tab. I. The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [45] and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume. The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
IV Method
IV-A Overview
@@
IV-B Tri-Plane Selection
Rather than videos, 3D VEM images can be interpreted in different planes. In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically. This phenomenon can be further confirmed by the experiment shown in Table III in which tracking along the $y$ axis is more effective than tracking along the $x$ axis. Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed. Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig. 4a, xy-plane). In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape. Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$. TriSAM terminates when no plane is selected.
IV-C SAM-based Tracking
Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis. If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt. If we use the bounding box of the current segment as the box prompt, SAM may not track in the long term as the output enclosed in the bounding box is getting smaller and smaller. Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results. TriSAM terminates when SAM returns no prediction for the next image slice.
IV-D Recursive Seed Sampling
@@
To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels. Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints. Then, we add these points to the global seed tree whose root note is the input seed. To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction. For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions. During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction. With such a seed tree data structure, we run TriSAM recursively until all potential points are visited. For each endpoint, we re-run the above two TriSAM modules to track segments along directions that are different from the current tracking axis.
@@
Initialize the seed tree $SeedTree=s$
@@
V Experiments
In this section, we present TriSAM’s experimental results. We begin with settings and implementation details. Then we show the benchmark results. Finally, we perform ablations for further analysis.
V-A Experimental Settings
@@
TABLE II: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the significant amount of our proofreading effort.
@@
SAM+IoU tracking zero-shot 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
@@ V-B Benchmark Results
Results Analysis. The results are shown in Tab. II. First, both the Color Thresholding and SAM+IoU Tracking methods exhibit significant performance variability across three volumes, highlighting the diversity of our dataset and the sensitivity of these methods to different species. Furthermore, both of these unsupervised methods demonstrate relatively poor performance, underscoring the challenges of the zero-shot setting in the BvEM dataset. Additionally, the 3D UNet, as a supervised learning approach, also yields subpar results, indicating poor generalization of models trained with limited data. Finally, TriSAM significantly outperforms other methods as it not only accurately segments the boundary but also tracks the blood vessels in the long term.

### v3

@@ Abstract
While imaging techniques at macro and mesoscales have garnered substantial attention and resources, microscale Volume Electron Microscopy (vEM) imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure. In this paper, we address a significant gap in this field of neuroimaging by introducing the first-in-class public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in vEM images. Our BvEM benchmark is based on vEM image volumes from three mammals: adult mouse, macaque, and human. We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation. Furthermore, we developed a zero-shot cortical blood vessel segmentation method named TriSAM, which leverages the powerful segmentation model SAM for 3D segmentation. To extend SAM from 2D to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points. This approach effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning. Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species. Our dataset, code, and model are available online at https://jia-wan.github.io/bvem.
{IEEEkeywords}
@@
Figure 1: Comparison of imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details. (b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all types of cells with details.
1 Introduction
With around 2% of body weight, our brain receives around 20% of blood supply. Most of the energy and nutrients are consumed by the neurons, and neuron function is sensitive to the blood supply [3, 4]. The blood supply can even be adjusted following the consumption of brain regions, called neurovascular coupling [5]. Alterations of blood vessel structures are observed in many brain diseases, e.g., Alzheimer’s and vascular dementia [6]. Thus, blood vessels in the brain have been extensively investigated with various imaging modalities at different resolutions (Figure 1a). Compared to the macro-level imaging (e.g., CT [1] , MRI [7]) and mesoscale-level imaging (e.g., light microscopy [2]), volume electron microscopy (VEM) [8] can further reveal the detailed ultrastructure including all vascular cells (Figure 1b) for in-depth analysis. However, no large-scale annotated VEM dataset exists to develop and evaluate automated 3D blood vessel segmentation methods.
@@ 1 Introduction
Thus, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images to foster segmentation method development. The raw image volumes are from recent publications, which are the largest for each of the three mammals: mouse, macaque, and human acquired at different VEM facilities. We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness. However, on the BvEM dataset, existing 3D blood vessel segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the imaging pipeline and the complexity of the 3D blood vessel morphology. Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature.
@@
To address these challenges, we propose a zero-shot 3D segmentation method, TriSAM, leveraging the recent 2D segmentation foundation model, Segment Anything Model (SAM) [15], to handle the appearance diversity. To re-purpose SAM for 3D segmentation, we developed a multi-seed segment tracking framework based upon the video object segmentation method with SAM [16]. As it is easier to track 2D blood vessel segments along the blood flow direction, TriSAM selects the best 2D plane for SAM-based tracking and further introduces a recursive seed sampling strategy that performs tri-plane selection at the potential turning points for efficiency. The proposed method is similar to Flood-Filling Networks (FFN) [17] by extending the mask in 3D space. Instead of training a 3D neural network, we utilize SAM for prediction since it generalize better. Second, FFN densely extends to 3D while the proposed method extends along the blood vessels. Finally, FFN uses only one seed while we are using growing seeds for long-term tracking. Experimental results demonstrate that the proposed TriSAM method significantly outperforms the prior state-of-the-art methods on the proposed BvEM benchmark across all three species.
2 Related Work
@@
Blood vessel segmentation methods. Most existing VEM image segmentation algorithms were developed for neurons [18] and synapses [19, 20, 21]. The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11]. This marks a significant milestone in connectomics research. The blood vessels were segmented manually in the human cortex [12] and automatically using fully convolutional neural network (FCN) [22] or 3D U-Net [11, 23, 24]. Due to the burden of annotation, efforts have been made to decrease the need for annotations [25]. For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10]. Briefly, the blood vessels could be enhanced based on the tubular nature using filters like diffusion and ridge filtering. Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
@@
Segment anything-based methods. As a foundation model for image segmentation, the recently proposed Segment Anything Model [15] has garnered significant attention and has been extended to a variety of domains [26, 27, 28, 29] including object tracking [26, 16], image inpainting [27], image mattting [30], super-resolution [31], 3D point cloud [32], and image editting [33]. Despite SAM’s remarkable generalization capabilities, it still encounters some challenges in practical applications, One of these challenges is the huge computation costs due to the heavyweight image encoder. FastSAM [34] adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed. MobileSAM [35] proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size. Therefore the MobileSAM is employed in our proposed method. One important challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical [36] or biological [37] images, particularly in the context of 3D data. Although SAM demonstrates impressive generalization capabilities, it faces challenges in specialized domains like medical [36, 38] or biological images [37], especially in handling 3D data. [29] assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties. [39] extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability. To address the domain gap between natural and medical images, SAM-Adapter [36], SAM-Med2D [40], and Medical SAM Adapter [41] introduced Adapter modules and trained the Adapter with medical images. They attained good performance on various medical image segmentation tasks. MedSAM [28] adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results. MicroSAM [37] also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data. Unlike these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
@@
3 BvEM Dataset
3.1 Dataset Description
Sample information. We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [11], superior temporal gyrus from an adult macaque [42], and temporal lobe from an adult human [12]. Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
Table 1: Dataset information. Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined. The last column shows the max/total length of the blood vessel instances.
@@ 3.1 Dataset Description
Resolution standardization. We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and the nature of capillary diameter. Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content. As shown in Fig. 2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
3.2 Dataset Annotation
@@
Semi-automatic segmentation error correction. We used the VAST lite software [43] which can accelerate the proofreading process by using provided segmentation results as the drawing or filling template instead of delineating segments manually.
Quality control. We used 3D visualizations and skeleton analysis to detect remaining segmentation errors. Each volume was proofread by two neuroscience experts in multiple rounds until no disagreement.
@@
3.3 Dataset Statistics
We summarize the dataset details in Tab. 1. The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [44] and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume. The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
@@ 3.3 Dataset Statistics
4 Method
4.1 Overview
@@
4.2 Tri-Plane Selection
Rather than videos, 3D VEM images can be interpreted in different planes. In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically. This phenomenon can be further confirmed by the experiment shown in Table 3 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis. Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed. Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig. 5a, xy-plane). In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape. Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$ (Select in Alg. 1). TriSAM terminates when no plane is selected.
4.3 SAM-based Tracking
Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis (SAMBasedTrack in Alg. 1). If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt. If we use the bounding box of the current segment as the box prompt, SAM may not track in the long term as the output enclosed in the bounding box is getting smaller and smaller. Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results. TriSAM terminates when SAM returns no prediction for the next image slice (probability $<\tau$ ).
4.4 Recursive Seed Sampling
@@
To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels (RecursiveSample in Alg. 1). Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints. Then, we add these points to the global seed tree whose root note is the input seed (LeafSplit in Alg. 1). To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction. For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions (Valid in Alg. 1). During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction. With such a seed tree data structure, we run TriSAM recursively until all potential points are visited. For each endpoint, we re-run the above two TriSAM modules to track segments along directions that are different from the current tracking axis.
@@
Initialize the seed tree $SeedTree=\{s\}$
@@
5 Experiments
5.1 Experimental Settings
@@
Table 2: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the significant amount of our proofreading effort.
@@
SAM+IoU tracking[15] zero-shot 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
@@ 5.2 Benchmark Results
Results Analysis. The results are shown in Tab. 2. First, both the Color Thresholding and SAM+IoU Tracking methods exhibit significant performance variability across three volumes, highlighting the diversity of our dataset and the sensitivity of these methods to different species. Furthermore, both of these unsupervised methods demonstrate relatively poor performance, underscoring the challenges of the zero-shot setting in the BvEM dataset. Additionally, the 3D UNet, as a supervised learning approach, also yields subpar results, indicating poor generalization of models trained with limited data. Finally, TriSAM significantly outperforms other methods as it not only accurately segments the boundary but also tracks the blood vessels in the long term.